# Melding the Serverless Control Plane with the Conventional Cluster Manager for Speed and Resource Efficiency

### v3

@@ 6. Evaluation
6.1. PulseNet Sensitivity Studies
@@
As discussed in §4.3, PulseNet’s performance is primarily affected by two parameters: (1) keepalive duration and (2) metric filtering percentile. Below, we study PulseNet sensitivity to these parameters and identify the sweet spot in the performance-cost trade-off to choose these parameters’ values when running sampled production traces.
6.1.1. Keepalive Duration.
@@
The keepalive duration determines how long a Regular Instance can remain idle before it is terminated. A smaller keepalive increases the cold start probability, degrading performance. A higher keepalive causes resource wastage, increasing cost. We sweep the keepalive duration from 2s to 600s to measure its impact on system performance and cost. As Figure 7(a) shows, when the keepalive duration reaches 60s, further increasing the keepalive duration significantly increases cost with minimal performance improvement. Thus, we set the keepalive duration to 60s.
6.1.2. Filtering Percentile.
@@
The filtering percentile determines the confidence threshold to create new function instances. Specifically, a lower filtering percentile makes the system more likely to create new instances, potentially leading to resource over-allocation and higher costs. A higher filtering percentile makes the system more conservative in creating new instances, potentially increasing the number of cold starts and degrading performance. We vary the percentile from 25% to 99%. Figure 7(b) shows that PulseNet achieves the best balance between performance and cost when the filtering percentile is 50%, which we use below.
@@
Given the obtained results, we choose the 60-second keepalive period and the 50-th percentile as metric filtering percentile in PulseNet as our parameters for evaluation.
6.2. Control Plane Performance Analysis
@@
In this section, we evaluate PulseNet’s control plane performance from three perspectives: instance creation delays, scheduling delays, and the system’s sensitivity to instance creation delays.
6.2.1. Instance Creation Delay Breakdown
@@
We first compare Regular and Emergency Instance creation delays (Figure 8). PulseNet eliminates all interactions with the cluster manager and uses a pool of pre-allocated local IP addresses to create the Emergency Instance in under 150ms, nearly 10 faster than creating a Regular Instance. Importantly, Regular Instance creation delays always occur off the critical path when handling invocations in PulseNet. Hence, PulseNet exposes only the cold-start delays of Emergency Instances to the application.
6.2.2. Scheduling Delays Analysis
@@
We continue the evaluation with the scheduling delays of different serverless systems. Scheduling delay includes cold start time, data plane queuing, request routing, and load balancing. To measure the scheduling delay, we subtract the function’s execution time from the end-to-end invocation latency. Figure 9 shows the distribution of average scheduling delay for each system.
@@
As shown in Figure 9, Knative, Dirigent, and PulseNet show median delays of 1s, 200ms, and 150ms, respectively, matching their instance creation times. In addition to faster instance creation delays, PulseNet eliminates the worst-case latencies of up to 4 seconds, which occur sporadically for the functions running atop Knative and Dirigent due to their window-based autoscaling policies. Kn-NHITS and Kn-Sync, due to the high instance retention, have 40-50% of functions experiencing warm starts below 20ms. However, the remaining functions are delayed up to 2s due to slow instance creation. Kn-LR shows high latency, up to 4 seconds, indicating its high prediction error. Compared with the above baselines, PulseNet reduces worst-case scheduling delays by quickly creating the Emergency Instances while minimizing cluster resource usage.
6.2.3. Sensitivity to Various Instance Creation Delays
@@
Different providers use different deployment and sandbox technologies, from regular VMs to FaaS-specialized solutions, to isolate Regular Instances. Hence, we evaluate the impact of their corresponding instance creation delays on the performance of the baseline systems and PulseNet. We use KWOK (kwo, 2025) to set up the simulated instance creation delay from 100ms, as in container bootstrapping, to 100s, similar to the booting time of a general-purpose VM, and measure the performance of different serverless systems under the same sampled production trace.
@@
In summary, by creating Emergency Instances for excessive traffic, PulseNet effectively eliminates worst-case scheduling delays. Furthermore, based on this design, even in environments with high instance creation delays, PulseNet delivers more stable performance
6.3. Control Plane Resource Efficiency Analysis
@@
Next, we analyze PulseNet’s control plane resource efficiency from three perspectives: instance creation rate, as well as CPU and memory utilization across the cluster.
6.3.1. Instance Creation Rate
@@
In this experiment, we measure the instance creation rate over time while executing the Azure Functions trace and show the results in Figure 11(a). A lower instance creation rate reduces the control plane overhead and improves system stability.
@@
As shown in Figure 11(a), Kn-Sync has the lowest instance creation rate (0.1 instances/s) as it uses a long 10-minute keepalive duration, allowing most instances to be reused. Knative and Dirigent show comparable instance creation rates (1.8 and 1.6 instances/s, respectively) as they employ similar autoscaling policies to respond to current traffic. The prediction model-based Kn-NHITS and Kn-LR have the highest instance creation rates (2.4 and 11.3 instances/s, respectively). This is because they need to adjust the number of instances based on prediction results frequently. Compared with Knative, PulseNet reduces the instance creation rate by 60% to 0.8 instances/s. This is because PulseNet’s filtering mechanism intelligently avoids creating unnecessary Regular Instances for functions with little, instead handling those with Emergency Instances. This design can reduce the computational overhead within the control plane of the serverless system and improve system stability. We study the additional computational resources necessary to handle the instance creations in the following section.
6.3.2. Cluster-wide CPU Cycles Utilization Breakdown
@@
We define CPU overhead as additional CPU consumption beyond the resources needed to serve the user functions, which incur costs for the service providers. CPU overhead includes all control plane components (sandbox management, traffic predictions, health and metric gathering), data plane components (load balancer, ingress), and any computation overhead produced by prediction mechanisms.
As shown in Figure 11(b), Emergency Instances account for only 10% of total CPU usage by the instances. PulseNet has similar overhead as Knative (24%) since it adds extra computational overhead for Emergency Instance management, while reducing the load on the control plane due to lower instance creation rate.
Prediction-based systems, Kn-NHITS and Kn-LR, have the highest CPU overhead due to additional compute resources required for predictions and handling more instance creations (we exclude their training time as an overhead): 45% and 64%, respectively.
@@
Dirigent can handle the same instance creation rate as Knative with 2% computational overhead due to its lighter cluster management implementation. Kn-Sync, due to its longer keepalive period, reduces the frequency of instance creation, thereby decreasing CPU overhead to 9%. Although PulseNet CPU overhead is higher than that of Kn-Sync, PulseNet achieves a better balance of resource usage by trading slightly increased CPU overhead for significantly reduced memory usage. We will analyze this in depth in Section 6.3.3.
6.3.3. Memory Usage
In this section, we evaluate the memory usage of various serverless systems by replaying the Azure Functions trace.
We show our results in Figure 12.
Kn-Sync exhibits the highest memory usage because it employs a long keepalive period of 10 minutes to handle requests, sing 7 more memory for idle instances than for active ones.
The prediction accuracy of Kn-NHITS and Kn-LR leads to overprovisioning of instances of 5 and 4.5, respectively.
Dirigent, due to its fast instance creation speed, can create instances more quickly, reducing queuing and preventing subsequent overreaction by the autoscaler.
PulseNet uses our proposed metric filtering technique that allows some of the requests to be handled by an Emergency Instance without creating a long-living Regular Instance. These Emergency Instances contribute 10% of total non-idle instance memory usage. Compared with Knative and Kn-Sync, PulseNet improves memory utilization by 8% and 60%, respectively.
PulseNet uses 5% more memory for instances than Dirigent, which could be improved with better metric filtering mechanisms (§4.3).
@@ 6.3.3. Memory Usage
In summary, through our proposed dual-path control plane architecture and intelligent filtering mechanism, PulseNet achieves the best balance between memory and CPU resources among Kubernetes-compatible systems. Specifically, PulseNet reduces the memory utilization by 8-60% compared to the baselines (and only 5% more than the FaaS-specialized Dirigent) and imposes negligible CPU overhead.
6.4. Trade-off Space Analysis of Performance & Cost
@@ 6.4.1. Experiments with a Real System
In this section, we evaluate the trade-off between performance and cost for different systems. We obtain the performance-cost trade-off for different systems by varying instance retention-related parameters (e.g., keepalive period, autoscaling window) from 6 seconds to 10 minutes. As shown in Figure 13, PulseNet achieves the best tradeoff between performance and cost. Compared with Knative and its variants (such as Kn-NHITS and Kn-LR), PulseNet achieves higher performance at lower resource costs. Specifically, PulseNet outperforms Kubernetes-compatible systems with synchronous control planes by 1.5-3.5 at 8-70% lower cost, and surpasses asynchronous counterparts by 1.7-3.5 at 3-65% lower cost. PulseNet achieves 35% faster end-to-end performance at a comparable cost to the Dirigent system. The reason behind PulseNet’s optimal tradeoff is its dual-track control plane design. This design effectively manages sustainable traffic with Regular Instances and excessive traffic with Emergency Instances, thereby achieving higher performance with fewer resource costs.
6.4.2. Large-Scale Experiments
@@
Then, we use the same methodology and metrics as in Section 6.4.1 to evaluate the performance-cost tradeoffs in the large-scale cluster by simulating 50 worker nodes with KWOK and running real control plane components. The experimental results show that in this large-scale cluster, PulseNet still demonstrates significant improvements compared with state-of-the-art systems. Specifically, PulseNet outperforms Kubernetes-compatible systems with both synchronous and asynchronous control planes by up to 3 at up to 3 lower cost. As the cluster size increases, traditional systems encounter severe control plane congestion issues. Additionally, system configurations that performed well in smaller clusters can overload the control plane under larger workloads. PulseNet effectively addresses this problem through the filtering mechanism between its expedited and standard tracks.
7. Related Work
@@
Cluster manager designs. Both practitioners and academics explore many designs for datacenter cluster managers primarily targeting long-running services and microservices (k8s, 2025a; Hindman et al., 2011; Vavilapalli et al., 2013; Schwarzkopf et al., 2013; Isard et al., 2009; Gog et al., 2016). Some works (Ousterhout et al., 2013; Schwarzkopf et al., 2013) explore trade-offs between centralized and decentralized cluster manager designs. Other works (Delimitrou and Kozyrakis, 2013; Delimitrou et al., 2015; Delimitrou and Kozyrakis, 2014) study the interference among the services co-located in the same cluster. The advent of Borg (Verma et al., 2015; Tirmazi et al., 2020) has mostly closed the research area for general-purpose cluster managers, making its open-source alternative Kubernetes (k8s, 2025a) the system of choice for long-running services and the foundational ecosystem for cloud offerings. PulseNet is fully compatible with the cluster managers, such as Kubernetes, benefiting from their mature ecosystem and optimizations. Although we show that conventional control planes based on Kubernetes deliver enough throughput to avoid queuing delays when creating instances at the datacenter scale, KOLE (Zhang et al., 2022) adapts Kubernetes for large-scale edge deployments but foregoes run-time instance creation essential for highly-dynamic serverless workloads.
@@
FaaS-Oriented Control Plane Optimizations. Recent works explore FaaS-specialized control planes, specifically focusing on FaaS workloads and often excluding the BaaS services from their consideration. Some works (Shahrad et al., 2020; Liu et al., 2024; Roy et al., 2022; Joosen et al., 2023; Mittal et al., 2021) focus on predicting the arrival time or request concurrency of future invocations to pre-allocate function instances ahead of time. Ilúvatar (Fuerst et al., 2023) reduces the scheduling overhead for warm starts. PulseNet is complementary to the above works, as it supports predictor-informed scaling in the standard track, and uses the conventional components in its warm-start path. Finally, other works (Cvetkovic et al., 2024; Singhvi et al., 2021; Mvondo et al., 2021; Fuerst and Sharma, 2021) take a radical approach designing FaaS-only scalable high-throughput control planes, foregoing the benefits of FaaS and BaaS application components colocation in the same cluster.
@@
BaaS-Oriented Control Plane Optimizations. Several works have designed BaaS-oriented control plane designs tailored for serverless applications. Some works (Romero et al., 2021; Klimovic et al., 2018; Mahgoub et al., 2021) optimize BaaS control planes for ephemeral storage by automatically scaling cache resources and dynamically selecting storage backends based on application workload patterns. Other works (Jesalpura et al., 2025; Yu et al., 2023; Du et al., 2022) propose fast technologies to facilitate data movement across functions, transparently to the application. PulseNet is complementary to these BaaS-specialized control planes, which can potentially be implemented as Kubernetes scheduler plug-ins.
8. Conclusion

### v4

@@ 7 Evaluation
7.1 PulseNet Sensitivity Studies
@@
As discussed in §4.3, PulseNet’s performance is primarily affected by two parameters: (1) keep-alive duration and (2) metric filtering percentile. Below, we study PulseNet sensitivity to these parameters and identify the sweet spot in the performance-cost trade-off to choose these parameters’ values when running sampled production traces.
@@
Keep-Alive Duration. The keep-alive duration determines how long a Regular Instance can remain idle before it is terminated. A smaller keep-alive increases the cold start probability, degrading performance. A higher keep-alive causes resource wastage, increasing cost. We sweep the keep-alive duration from 2s to 600s to measure its impact on system performance and cost. As Figure 7(a) shows, when the keep-alive duration reaches 60s, further increasing it significantly increases the cost with minimal performance improvement. Thus, we set the keep-alive duration to 60s.
@@
Filtering Percentile. The filtering percentile determines the confidence threshold to create new function instances. Specifically, a lower filtering percentile increases the likelihood of creating new instances, leading to resource over-allocation and higher costs. A higher filtering percentile makes the system more conservative in creating new instances, increasing the number of cold starts and degrading performance. We vary the percentile from 25% to 99%. Figure 7(b) shows that PulseNet achieves the best balance between performance and cost when the filtering percentile is 50%, which we use below.
@@
Given the obtained results, we choose the 60-second keep-alive period and the 50-th percentile as metric filtering percentile in PulseNet as our parameters for evaluation.
7.2 Control-Plane Performance Analysis
@@
In this section, we evaluate PulseNet’s control-plane performance from three perspectives: instance creation delays, scheduling delays, and sensitivity to instance creation delays.
7.2.1 Instance Creation Delay Breakdown
@@
We first compare the creation delays of Regular and Emergency instances. PulseNet eliminates all interactions with the cluster manager and uses a pool of pre-allocated local IP addresses to create the Emergency Instance in under 150ms, nearly 10 faster than creating a Regular Instance in Knative (Figure 4). Importantly, Regular Instance creation delays always occur off the critical path when handling invocations in PulseNet. Hence, PulseNet exposes only the cold-start delays of Emergency Instances to the application.
7.2.2 Scheduling Delays Analysis
@@
We continue the evaluation with the scheduling delays of different serverless systems. Scheduling delay includes cold-start time, data-plane queuing, request routing, and load balancing. To measure the scheduling delay, we subtract the function’s execution time from the end-to-end invocation latency. Figure 8 shows the distribution of average scheduling delay for each system.
@@
As shown in Figure 8, Knative, Dirigent, and PulseNet show median delays of 1s, 200ms, and 150ms, respectively, matching their instance creation times. In addition to faster instance creation delays, PulseNet eliminates the worst-case latencies of up to 4 seconds that sporadically occur in Knative and Dirigent due to their window-based autoscaling policies. Kn-NHITS and Kn-Sync, due to their high instance retention, have 40-50% of functions that experience warm starts below 20ms. However, the remaining functions are delayed up to 2s due to slow instance creation. Kn-LR shows high latency, up to 4 seconds, indicating its high prediction error. Compared with the above baselines, PulseNet reduces worst-case scheduling delays by quickly creating Emergency Instances when no Regular Instances are present or when all Regular Instances are busy, while minimizing cluster resource usage.
7.2.3 Sensitivity to Various Instance Creation Delays
@@
Different providers use different deployment and sandbox technologies, from regular VMs to FaaS-specialized solutions, to isolate Regular Instances. Hence, we evaluate the impact of their corresponding instance creation delays on the performance of the baseline systems and PulseNet. We use KWOK [kwok] to set up the simulated instance creation delay from 100ms, as in container bootstrapping, to 100s, similar to the booting time of a general-purpose VM, and measure the performance of different serverless systems under the same sampled production trace.
@@
In summary, by creating Emergency Instances for excessive traffic, PulseNet effectively eliminates worst-case scheduling delays. Furthermore, based on this design, even in environments with high instance creation delays, PulseNet delivers more stable performance.
7.3 Control-Plane Resource Efficiency
@@
Next, we analyze PulseNet’s control-plane resource efficiency from three perspectives: instance creation rate and CPU and memory utilization across the cluster.
7.3.1 Instance Creation Rate
@@
In this experiment, we measure the instance creation rate over time while executing the Azure Functions trace and show the results in Figure 10(a). A lower instance creation rate reduces the control-plane overhead and improves system stability.
@@
As shown in Figure 10(a), Kn-Sync has the lowest instance creation rate (0.1 instances/s) as it uses a long 10-minute keep-alive duration, allowing most instances to be reused. Knative and Dirigent show comparable instance creation rates (1.8 and 1.6 instances/s, respectively) as they employ similar autoscaling policies. The prediction model-based Kn-NHITS and Kn-LR have the highest instance creation rates (2.4 and 11.3 instances/s, respectively). This is because they need to adjust the number of instances based on prediction results frequently. Compared with Knative, PulseNet reduces the instance creation rate by 60% to 0.8 instances/s. This is because PulseNet’s filtering mechanism intelligently avoids creating unnecessary Regular Instances for functions with little, instead handling those with Emergency Instances. This design can reduce the computational overhead in the serverless system’s control plane and improve system stability. We study the additional computational resources required to handle instance creation in the following section.
7.3.2 Cluster-wide CPU Cycles Utilization Breakdown
@@
We define CPU overhead as additional CPU consumption beyond the resources required to serve user functions, which incurs costs for service providers. CPU overhead includes all control-plane components (sandbox management, traffic predictions, health and metric gathering), data-plane components (load balancer, ingress), and any computation overhead produced by prediction mechanisms.
@@
As shown in Figure 10(b), Emergency Instances account for only 10% of total CPU usage by the instances. PulseNet has similar overhead as Knative (24%) since it adds extra overhead for Emergency Instance management, while reducing the load on the control plane due to fewer instances created. Prediction-based systems, Kn-NHITS and Kn-LR, have the highest CPU overhead due to additional compute resources required for predictions and handling more instance creations (we exclude their training time as an overhead): 45% and 64%, respectively. Dirigent can handle the same instance creation rate as Knative with 2% computational overhead due to its lighter cluster management implementation. Kn-Sync, due to its longer keep-alive period, reduces the frequency of instance creation, thereby decreasing CPU overhead to 9%. Although PulseNet CPU overhead is higher than that of Kn-Sync, PulseNet achieves a better balance of resource usage by trading slightly increased CPU overhead for significantly reduced memory usage. We will analyze this in depth in §7.3.3.
7.3.3 Memory Usage
In this section, we evaluate the memory usage of various serverless systems by replaying the Azure Functions trace. We show our results in Figure 11. Kn-Sync exhibits the highest memory usage because it employs a long keep-alive period of 10 minutes to handle requests, using 7 more memory for idle instances than for active ones. The prediction accuracy of Kn-NHITS and Kn-LR leads to overprovisioning of instances of 5 and 4.5, respectively. Dirigent, due to its fast instance creation speed, can create instances more quickly, reducing queuing and preventing subsequent overreaction by the autoscaler. PulseNet uses our proposed metric-filtering technique, which allows some requests to be handled by an Emergency Instance without creating a long-lived Regular Instance. These Emergency Instances account for 10% of the total non-idle instance memory usage. Compared with Knative and Kn-Sync, PulseNet improves memory utilization by 8% and 60%, respectively. PulseNet uses 5% more memory for instances than Dirigent, which could be improved with better metric filtering mechanisms (§4.3).
@@ 7.3.3 Memory Usage
In summary, with our proposed dual-path control-plane architecture and intelligent filtering mechanism, PulseNet achieves the best balance between memory and CPU resources among Kubernetes-compatible systems. Specifically, PulseNet reduces the memory utilization by 8-60% compared to the baselines (and only 5% more than the FaaS-specialized Dirigent) and imposes negligible CPU overhead.
7.4 Performance & Cost Trade-off Analysis
@@
Experiments with a Real System. In this section, we evaluate the trade-off between performance and cost for different systems. We obtain the performance-cost trade-off by varying instance retention-related parameters (e.g., keep-alive period, autoscaling window) from 6 seconds to 10 minutes. As shown in Figure 12, PulseNet achieves the best tradeoff between performance and cost. Compared with Knative and its variants (such as Kn-NHITS and Kn-LR), PulseNet achieves higher performance at lower resource costs. Specifically, PulseNet outperforms Kubernetes-compatible systems with synchronous control planes by 1.5-3.5 at 8-70% lower cost, and surpasses asynchronous counterparts by 1.7-3.5 at 3-65% lower cost. PulseNet achieves 35% faster end-to-end performance at a comparable cost to the Dirigent system. The reason behind PulseNet’s optimal tradeoff is its dual-track control-plane design. This design effectively manages sustainable traffic with Regular Instances and excessive traffic with Emergency Instances, thereby achieving higher performance with fewer resource costs.
@@
Large-Scale Experiments. Then, we use the same methodology and metrics to evaluate the performance-cost trade-offs in the large-scale cluster by simulating 50 worker nodes with KWOK and running real control-plane components. The experimental results show that at this larger scale, PulseNet still demonstrates significant improvements compared with state-of-the-art systems. Specifically, PulseNet outperforms Kubernetes-compatible systems with both synchronous and asynchronous control planes by up to 3 at up to 3 lower cost. As the cluster size increases, traditional systems experience severe control-plane congestion. Additionally, system configurations that performed well in smaller clusters can overload the control plane under larger workloads. PulseNet effectively addresses this problem through the filtering mechanism between its expedited and standard tracks.
7.5 Generalizability of PulseNet
Sensitivity to Per-Instance Concurrency. As some providers, such as Google Cloud Run [gcr], recommend setting up per-instance concurrency (CC) larger than 1, we compare PulseNet performance and cost to the Knative baseline with larger CC. For the Knative baseline, increasing CC from 1 to 10 improves performance by about 20%, with the speedup saturating at CC of 5. Hence, a higher CC value can modestly shrink the gap in performance between PulseNet and the baseline (as CC does not affect the expedited track of PulseNet, PulseNet performance does not change), but the latter remains more than 3 faster. Furthermore, operating Knative at CC of 5 increases the overall operational cost by 3, assuming that concurrency-enabled instances incur proportionally higher resource costs.
Real-World Applications.
We verify the generality of our findings by running real-world workloads from vSwarm written in Go, Python, and Node.js. We observe a 4.3 reduction in tail slowdowns, as shown in Figure 13, with a negligible impact on the number of instances in PulseNet relative to the Knative baseline. This performance gain is similar to previous experiments, however, with a larger performance improvements over the baseline with real functions compared to with synthetic ones, confirming the effectiveness of PulseNet in real-world deployments.888We attribute those differences to the limitations of mapping 400 functions from Azure Trace to a set of replicated 11 vSwarm functions.
8 Related Work
@@
Cluster manager designs. Both practitioners and academics explore many designs for datacenter cluster managers primarily targeting long-running services and microservices [k8s, hindman:mesos, vavilapalli:yarn, schwarzkopf:omega, isard:quincy, gog:firmament]. Some works [ousterhout:sparrow, schwarzkopf:omega] explore trade-offs between centralized and decentralized cluster manager designs. Other works [delimitrou:paragon, delimitrou:tarcil, delimitrou:quasar] study the interference among the services co-located in the same cluster. The advent of Borg [verma:borg, tirmazi:borg] has mostly closed the research area for general-purpose cluster managers, making its open-source alternative Kubernetes [k8s] the system of choice for long-running services and the foundational ecosystem for cloud offerings. PulseNet is fully compatible with the cluster managers, such as Kubernetes, benefiting from their mature ecosystem and optimizations. Although we show that conventional control planes based on Kubernetes deliver enough throughput to avoid queuing delays when creating instances at the datacenter scale, KOLE [zhang:kole] adapts Kubernetes for large-scale edge deployments but foregoes run-time instance creation essential for highly-dynamic serverless workloads.
@@
FaaS-Oriented Control-Plane Optimizations. Recent works explore FaaS-specialized control planes, specifically focusing on FaaS workloads and often excluding the BaaS services from their consideration. Some works [shahrad:serverless, liu:jiagu, roy:icebreaker, joosen:how, mittal:mu] focus on predicting the arrival time or request concurrency of future invocations to pre-allocate function instances ahead of time. Ilúvatar [fuerst:iluvatar] reduces the scheduling overhead for warm starts. PulseNet is complementary to the above works, as it supports predictor-informed scaling in the standard track, and uses the conventional components in its warm-start path. Finally, other works [cvetkovic:dirigent, singhvi:atoll, mvondo:ofc, fuerst:faascache] take a radical approach designing FaaS-only scalable high-throughput control planes, foregoing the benefits of FaaS and BaaS application components colocation in the same cluster.
@@
BaaS-Oriented Control-Plane Optimizations. Several works have proposed BaaS-oriented control-plane designs tailored for serverless applications. Some works [romero:faast, klimovic:pocket, mahgoub:sonic] optimize BaaS control planes for ephemeral storage by automatically scaling cache resources and dynamically selecting storage backends based on application workload patterns. Other works [Jesalpura:Shattering, yu:following, du:molecule] propose fast technologies to facilitate data movement across functions, transparently to the application. PulseNet is complementary to these BaaS-specialized control planes, which can potentially be implemented as Kubernetes scheduler plug-ins. Furthermore, SPRIGHT [qi:spright] and Nightcore [jia:nightcore] mitigate overheads in data movement between co-located instances by replacing TCP/IP networking with eBPF and IPC, respectively. FUYAO [liu:fuyao] extends this approach to optimize intra-node communication by eliminating the TCP/IP stack, in addition to DPU-accelerated inter-node data movement. Similarly, Apiary [kraft:apiary] places the function execution into the database’s user-defined functions. These works underscore the benefits of data locality, which can be achieved through scheduling policies in conventional cluster managers such as Kubernetes [bazzaz:preventing, zhang:faster].
Other optimizations for serverless systems: cold-start delays, microvms, and hypervisors. Many recent works [szekely:sigmaos, ruan:quicksand, kuchler:dandelion, segarra:granny, fried:junction, huang:trenv] redesign the runtimes to improve the performance and efficiency of serverless and cloud deployments. Other works [fu:serverlessllm, zeng:medusa] apply similar approaches for GPU-centric workloads. These works are orthogonal to PulseNet and can be seamlessly incorporated into the PulseNet tracks.
9 Conclusion